40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Oct 2 23:15:02 2019, gautam, Update, LSC, FPMI locking FPMIlocked.png
    Reply  Thu Oct 3 11:01:18 2019, rana, Update, LSC, (PR)FPMI locking 
Message ID: 14926     Entry time: Wed Oct 2 23:15:02 2019     Reply to this: 14928
Author: gautam 
Type: Update 
Category: LSC 
Subject: FPMI locking 

Summary:

I was able to lock the FPMI. The lock was quite stable. However, the fluctuations in the ASDC power suggest that it will be difficult to make a DC measurement of the contrast defect in this configuration. This problem can be circumvented in part by some electronics tuning. However, the alignment jitter couples some HOM light which is an independent effect. Can this be a good testbed for the proposed AS WFS system? 

Details:

  • First, the arm cavities were locked and TRX/TRY were maximized using ASS.
  • Next, AS55_Q-->MICH_A (MICH-->BS) matrix element was set to 1 in the LSC input (output) matrix. The trigger was set to always on.
  • AS55 digital demod phase was -37 degrees.
  • I was then able to increase the gain on the MICH servo and turn on some integrators without any problem.
  • Some guesswork had to be done to get the correct sign. Final servo gain used was -0.8. 

I didn't do any serious budgeting yet - need to think about / do some modeling on how this configuration can be made useful.

Attachment 1: FPMIlocked.png  43 kB  Uploaded Thu Oct 3 00:20:02 2019  | Hide | Hide all
FPMIlocked.png
ELOG V3.1.3-